Your firm purchases goods from its supplier on terms of 2/10, Net 45. What is the effective annual cost to your firm if it chooses not to take advantage of the trade discount offered?
Answer #1

The firm will pay either 98 on or before 10 days or 100 with in 45 days. The difference in days 45-10=35days.

Effective annual cost = (100/98)^(365/35)-1 = 23.45 %